Session: Feb 1999 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Fri Feb 26 11:24:15 MET 1999) Success
No known problems.

Noto (C.Trigilio at Fri Feb 26 11:29:54 MET 1999) Minor failures
BBC12 not locked to the right frequency till 6:30 UT. No other problem.

Onsala (A. Polatidis at Fri Feb 26 14:19:15 MET 1999) Success

Westerbork (Tony Foley at Fri Feb 26 22:39:23 MET 1999) Success
No known problems

Urumqi (Liu Xiang at Mon Mar 1 09:27:03 MET 1999) Minor failures
The experiment started at UT 06:22:00, because of receiver problem. The parity sometimes is high.

Shanghai (Huang Xinyong at Mon Mar 1 11:56:12 MET 1999) Success
No problems during the observation. It is raining.

Jodrell (Alastair Gunn at Tue Mar 2 05:40:30 MET 1999) Minor failures
Telescope still slewing at start of experiment, on source at 06:10 UT. No other known problems.

Cambridge (Alastair Gunn at Tue Mar 2 05:41:05 MET 1999) Failure
Headstack position incorrect throughout entire experiment due to blown HV fuse in inchworm driver unit.

Effelsberg (dgraham at Tue Mar 2 13:38:07 MET 1999) Success

Calibration (Jean\-Francois Desmurs at Wed Mar 3 18:34:12 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: c99l1.antab. Read the comments in the file to get further information.